### v1

# Kramers nodal lines in intercalated TaS2 superconductors

Abstract
Abstract
Kramers degeneracy is one fundamental embodiment of the quantum mechanical nature of particles with half-integer spin under time reversal symmetry. Under the chiral and noncentrosymmetric achiral crystalline symmetries, Kramers degeneracy emerges respectively as topological quasiparticles of Weyl fermions and Kramers nodal lines, anchoring the Berry phase-related physics of electrons. However, an experimental demonstration for ideal Kramers nodal lines (KNLs) well isolated at the Fermi level is lacking. Here, we establish a class of noncentrosymmetric achiral intercalated transition metal dichalcogenide (TMD) superconductors with large Ising-type spin-orbit coupling, represented by InxTaS2, to host an ideal KNL phase. We provide evidence from angle-resolved photoemission spectroscopy (ARPES) with spin resolution, angle-dependent quantum oscillation (QO) measurements, and ab-initio calculations. Our work not only provides a realistic platform for realizing and tuning KNLs in layered materials, but also paves the way for exploring the interplay between KNLs and superconductivity, as well as applications pertaining to spintronics, valleytronics, and nonlinear transport.
{affiliations}
Department of Physics and Astronomy, Rice Center for Quantum Materials (RCQM), Rice University, Houston, Texas 77005, USA
New Technologies Research Center, University of West Bohemia, Plzen 301 00, Czech Republic
Department of Physics, Hong Kong University of Science and Technology, Clear Water Bay, Hong Kong, China
Department of Physics, University of California, Berkeley, California 94720, USA
National High Magnetic Field Laboratory, Tallahassee, Florida 32310, USA
Stewart Blusson Quantum Matter Institute, University of British Columbia, Vancouver, British Columbia V6T 1Z4, Canada
Advanced Light Source, Lawrence Berkeley National Laboratory, Berkeley, California 94720, USA
Stanford Synchrotron Radiation Lightsource, SLAC National Accelerator Laboratory, 2575 Sand Hill Road, Menlo Park, California 94025, USA
National Synchrotron Light Source II, Brookhaven National Lab, Upton, NY 11973, USA
Department of Physics & Astronomy, University of British Columbia, Vancouver, British Columbia V6T 1Z1, Canada
Canadian Institute for Advanced Research, Toronto, Ontario M5G 1M1, Canada
Materials Science Division, Lawrence Berkeley National Laboratory, Berkeley, California 94720, USA
Department of Physics and Texas Center for Superconductivity, University of Houston, Houston, Texas 77204, USA
Introduction
Symmetry plays a ubiquitous role in dictating the electronic properties of solids, enriched by the introduction of topology into the field of condensed matter [1]. In particular, recent developments have recognized the presence of topological degeneracies in the electronic band structure originated from crystalline symmetries, such as non-symmorphic [2, 3], chiral [4, 5], and achiral [6] operations, which could be further intertwined with magnetic [7, 8, 9] and charge order [10]. In the scenario of non-symmorphic symmetry, Dirac- and Weyl-type band crossings [11], unconventional multi-fold fermions[12, 13, 14], and hourglass fermions [15, 16, 17] could arise from the glide-mirror or screw-axis symmetries. Meanwhile, chiral and noncentrosymmetric achiral little group symmetries emphasize the absence of inversion, albeit bearing overlapping space groups with the non-symmorphic ones. Crucially, mirror or roto-inversion symmetries must be present (absent) in the achiral (chiral) structure. This difference is the key in determining the distinction between Kramers-Weyl fermions pinned at time reversal invariant momenta (TRIM) in chiral crystals [4] and the type-\@slowromancapi@ Kramers nodal lines (KNLs) connecting TRIM in noncentrosymmetric achiral crystals [6]. These symmetries, if bearing topological quasiparticles located near the Fermi level and isolated from trivial bands, could generate distinct transport, thermal, and optical phenomena such as Berry curvature related anomalous Hall [18] and Nernst effects [19], chiral anomaly [20], dissipationless edge current [21], and circular photogalvanic effect [22]. Recently, chiral crystals have garnered renewed interest owing to their monopole-like orbital augular momentum texture [23], leading to promising aspects on orbital magnetotransport, while noncentrosymmetric achiral crystals hosting KNLs, a type of Dirac solenoid concentrating quantized Berry flux of $\pi$, still require identification of ideal material candidates and unequivocal experimental demonstration.
Since the theory prediction of KNLs [6], experimental work has suggested a few material platforms. These include time reversal symmetry breaking superconductors of the $T$RuSi ($T$ = Ti, Ta, Nb, and Hf) [24] and the LaPtSi [25] family, the paramagnetic state of SmAlSi [26, 27], and charge density wave (CDW)-driven KNLs in rare-earth tritellurides [28, 29]. However, in the case of rare-earth tritellurides, the approximated polar supercell derived from the x-ray diffraction (XRD) resolved incommensurate CDW superspace group [30, 28] directly contradicts with the observation of preserved inversion symmetry in the combined studies of Raman spectroscopy, rotational-anisotropy second harmonic generation, and other experimental techniques [31], hence casting doubt on the existence of KNLs driven by the CDWs in the series of systems. Moreover, these materials are all either multi-band systems near the Fermi level, $E_{\rm F}$, or have the KNLs located far away from $E_{\rm F}$.
Here, through comprehensive physical and thermodynamic properties characterization, angle-resolved photoemission spectroscopy (ARPES), and spin-resolved ARPES measurements, ab-initio calculations, and quantum oscillation measurements, we introduce a material class that exhibits an ideal KNL metallic phase, in the form of a noncentrosymmetric achiral intercalated transition metal dichalcogenide (TMD) family in the space group of $P\bar{6}m2$ [32, 33]. These are exemplified by InxTaS2 (x=1/2, 2/3, 1) (Fig. 1), all of which also feature superconducting ground states [34, 35, 36]. Due to the environment of the TMD layers (Figs. 1(a-b)), the KNL band topology is also coupled with spin-valley degrees of freedom (Figs. 1(c-d)). Such spin texture of KNLs in the intercalated TMD system under the $D_{3h}$ point group can be clearly visualized by the low energy band structure only considering spin-orbit coupling (SOC) terms (Fig. 1(d)), of which the momentum-dependent relativistic pseudospin splitting pattern is reminiscent of the recently generalized concept of nonrelativistic altermagnetism [37, 38]. However, the fact that the directional relativistic splitting is locked with the crystal structure symmetry, unlike the nonrelativistic altermagnetic splitting associated with antiferromagnetism, enables experimental investigation without complications of domain alignment. Furthermore, our theoretical analysis on the generic ideal KNL model at the Fermi level reveals several potential magnetic field-induced effects, including anomalous Hall effect (AHE) and Majorana zero modes that are theoretically possible in the $s$-wave superconducting phase. Additionally, the noncentrosymmetric achiral $P\bar{6}m2$ InxTaS2 associates the symmetries of the KNL material family with recent experimental reports of giant anomalous nonlinear transport [39] and strain-induced superconducting diode effect [40] in the isostructural PbTaSe2 with magnetic field-odd dependent behaviour.
Results
0.1 Design of Kramers nodal line metals and characterization of the InxTaS2 family.
The first step in constructing a noncentrosymmetric achiral crystal structure is to break the inversion symmetry. 2H-TaS2 and other TMDs of the 2H phase have been long known to have a centrosymmetric structure [41]. However, due to the local broken inversion symmetry constrained within a single layer, layer-resolved “hidden” spin polarization has been reported by surface sensitive ARPES techniques [42, 43]. In addition, extensive efforts focused on the monolayer and few-layer regime have shown evidence for unconventional Ising superconductivity that allows both spin-singlet and spin-triplet pairings [44, 45, 46, 47, 48, 49, 50, 51, 52]. To achieve inversion symmetry breaking in the bulk, we have intercalated 2H-TaS2 with trigonal indium layers, as shown in Figs. 1(a-b). The indium intercalation changes the 2H stacking to 1H stacking, and consequently, leads to broken inversion symmetry in the bulk of the InxTaS2 crystal structure. We also note that, as illustrated in Fig. 1(b), there is a structural difference between $x$ = 1/2, 2/3 and $x$ = 1: the In atoms are located above the Ta atoms for $x$ = 1/2 and 2/3 (left, Fig. 1b), while for $x$ = 1, the In atoms are above the S atoms (right, Fig. 1b). Such structural difference is supported by Rietveld refinements on the powder crystal XRD (see Supplementary Note 1). As a result, the Weyl nodal rings predicted in InTaS2 near $H$ point [35] (see Fig. 1(c) for the Brillouin zone (BZ) definition) based on a structure of In on top of Ta would instead become predicted Weyl points near $H$ [36]. Nonetheless, the KNL from the noncentrosymmetric achiral little group symmetry in InxTaS2 is fundamentally different from all the topological crossings predicted previously for similar compounds [35, 36, 53, 54]. Under the noncentrosymmetric achiral point group symmetry of InxTaS2, the electronic band structure from the two-band low energy Hamiltonian at $k_{z}$ = $\pi$ [6] only considering SOC terms shows an alternating pattern of pseudospin splitting, as shown in Fig. 1(d), while the black lines along $\Gamma-M$ indicate spin degeneracy and highlights the presence of the KNLs. The complete distribution of the KNLs is indicated by the purple lines in the first Brillouin zone (BZ) in Fig. 1(c). Furthermore, we visualize the realistic 3D Fermi surface (FS) of InTaS2 calculated from first-principles in the left panel of Fig. 1(e) showing the Kramers degeneracy along $\Gamma-M$ whenever the two FS sheets intersect. This can also be seen in the FS plotted for the $k_{z}$ = 0 plane where all Fermi pockets have pinched Kramers degeneracy when they cross the $\Gamma-M$ line (right of Fig. 1(e)).
[…]
0.2 Ideal Kramers nodal line metal In1/2TaS2 with spin-valley polarization.
To demonstrate the physics of KNL and its spin-orbital texture, a candidate material of ideal KNL metals is highly desired. Here, “ideal KNL” is defined as an isolated KNL band that crosses $E_{\rm F}$, with a large SOC that results in a band splitting away from the KNL momentum directions. Among the noncentrosymmetric achiral intercalated TMD material family, we identify In1/2TaS2 to exhibit such ideal properties. First, to demonstrate the KNL properties, we carried out ab-initio calculations on the bulk Bloch spectral functions of In1/2TaS2, where the coherent potential approximation was used on the In sites accounting for the random vacancy without extending to a CDW supercell. As outlined by the red box in Fig. 2(a), the calculation shows a single set of isolated bulk KNL bands along $M-\Gamma-A-L$ crossing the Fermi level. Away from this momentum path, the KNL band splits into two spin-orbital branches, as shown along $\Gamma-K-M$ and $A-H-L$, where the $k_{z}$ = $\pi$ dispersions show the most prominent splitting. Projecting the Green’s function with the $\sigma_{z}$ operator in Fig. 2(b), it is unambiguously shown that the spin is strictly degenerate along the KNLs and highly polarized in $s_{z}$ along the off-KNL directions due to the Ising SOC. The splitting is most clearly demonstrated along $A-H-L$, but hybridized with In orbitals along $K-M$. To directly observe the ideal KNL band topology and provide the smoking-gun evidence for the spin splitting, we performed spin-integrated ARPES (Fig. 2) and spin-resolved ARPES (Fig. 3) measurements on In1/2TaS2. As shown in Fig. 2(c), the In1/2TaS2 FSs consist of two concentric pockets around $\bar{K}$ and $\bar{K\textquoteright}$ and two hexagonal pockets centered around $\bar{\Gamma}$, with no clear evidence of CDW folding or gaps. The ARPES band dispersions along $\bar{\Gamma}-\bar{M}-\bar{K}-\bar{\Gamma}$ presented in Fig. 2(f) directly demonstrate the doubly-degenerate KNL crossing at $E_{\rm F}$ along $\bar{\Gamma}-\bar{M}$ and its clear splitting along $\bar{M}-\bar{K}-\bar{\Gamma}$. To theoretically capture the observed ARPES spectra with high accuracy, state-of-the-art one-step photoemission calculations for the FS and band dispersions along $\bar{\Gamma}-\bar{M}-\bar{K}-\bar{\Gamma}$ are carried out (Figs. 2(d) and (g)), showing an excellent agreement with ARPES data, confirming the validity of the random vacancy modelling on the In sites. Furthermore, the spin-projected one-step calculation indicates a spin-polarized FS encoding spin-valley degree of freedom at $\bar{K}$ and $\bar{K\textquoteright}$ (Fig. 2(e)).
The investigation of the spin-valley locking behavior associated with the KNLs using spin-resolved ARPES is presented in Fig. 3. As the spin polarization of the split bands should reverse on the two opposite directions away from the KNL, we can directly measure the spin polarization across a KNL. We carried out spin-resolved measurement of two pairs of momentum points on opposite sides of the $\Gamma-M$ KNL. $k_{1}$ and $k_{2}$ are along $\bar{K}-\bar{M}-\bar{K^{\prime}}$ (Fig. 3(a)). The spin up and down energy distribution curves (EDCs) selectively show the spin texture along the out-of-plane direction ($s_{z}$), and we plot them in direct comparison to the EDCs measured in spin-integrated mode in Fig. 3(b). Even though for both $k_{1}$ and $k_{2}$, the band closer to $E_{\rm F}$ exhibits weaker intensity as shown in the spin-integrated EDCs, the spin-resolved EDCs still show clear distinction for the split bands, namely for $k_{1}$ the lower binding energy band is dominantly spin up while the higher binding energy band is dominantly spin down. Furthermore, we can plot the $s_{z}$ polarization defined as $s_{z}=\frac{1}{S}\frac{I_{\uparrow}-I_{\downarrow}}{I_{\uparrow}+I_{\downarrow}}$ (Fig. 3(c), $S$ being the Sherman function, $I_{\uparrow}$ and $I_{\downarrow}$ the spectral intensity measured in the spin up and spin down channels, respectively), which clearly shows the opposite spin polarization for the two bands split from the KNL. $k_{2}$, on the opposite side of the $\Gamma-M$ KNL, shows reversed spin polarization of the two split bands (Figs. 3(d) - (e)). To demonstrate this spin polarization reversal even more clearly, we show similar measurements for a pair of momentum points slightly away from the $\bar{K}-\bar{M}-\bar{K^{\prime}}$ high symmetry direction where the photoemission matrix elements allow comparable intensity of the two split bands as observed on each spin-integrated EDC ($k_{3}$ and $k_{4}$) (Figs. 3(f)-(j)). The spin-resolved ARPES data ($s_{z}$) measured here shows an even better defined peak structure and showcases consistent spin texture with spin polarization reaching as high as nearly 80%. Therefore, the combined experimental and theoretical results definitively demonstrate the coupled spin-valley polarization in In1/2TaS2, associated with the KNL band topology and the underlying Ising-type SOC shown previously only for monolayer TMD materials, which could potentially be relevant to applications in spintronics and valleytronics.
0.3 Kramers nodal lines in the 3D electronic band structure of InTaS2.
As the noncentrosymmetric achiral symmetry is common across the family of intercalated TaS2, the KNL should be a universal property of all members of this family. In this section we present the evidence for KNLs for the stoichiometric InTaS2 in the $P\bar{6}m2$ space group. In addition, we also present the ARPES experimental measurement of the KNL band structure of the isostructural PbTaSe2 in the Supplementary Note 3. With full In intercalation, InTaS2 is more three-dimensional and exhibits stronger $k_{z}$ band dispersions when compared to In1/2TaS2 (Fig. 1(e)). Due to the $k_{z}$ broadening effect of the photoemission process, the measured FS of InTaS2 is more complex than the one of In1/2TaS2 (Fig. 4(a)). To account for potential surface states and the strong $k_{z}$ broadening effect in the vacuum ultraviolet ARPES spectra, we undertook ab-initio calculations projected to both In and S surface terminations. We find that the measured electronic structure can be best described by the sum of the calculations for both terminations. This is likely due to the fact that the In and S terminations have equal probabilities during the cleaving process. Therefore, the ARPES beam spot would probe the superposition of band structures under both terminations. The direct comparison of the measured and calculated FSs is shown in Figs. 4(a-b), where features such as the trefoil-like pattern centered at $\bar{K}$, the elliptical pockets centered at $\bar{M}$, and the $k_{z}$ broadened hexagonal pockets centered at $\bar{\Gamma}$ can be well reproduced. Importantly, we note that pairs of FSs are enforced to be degenerate whenever intersecting the $\Gamma-M$ direction, reflecting the properties of the KNL band topology. This can be further confirmed by the measured band dispersions along $\bar{\Gamma}-\bar{M}-\bar{K}-\bar{\Gamma}$ (Fig. 4(c)), where we also overlay the bulk band calculations for $k_{z}$ = 0 and $\pi$. Clearly, both the bulk calculations and the ARPES data demonstrate that the bands along $\Gamma-M$ are doubly degenerate KNLs and split into two spin-orbital branches along $M-K-\Gamma$. To better capture the band dispersions observed in Fig. 4(c) beyond bulk contributions, we also display the surface calculations projected onto the In surface, which show an excellent match with the ARPES data.
Since the FS of InTaS2 contains more 3D dispersions, as delineated in Fig. 1(e), and ARPES has significant spectral intensity contributions from $k_{z}$ broadening and surfaces, we use quantum oscillation (QO) measurements to provide a more complete investigation on the 3D bulk FS topology. The resistivity is measured as a function of the external magnetic field at different temperatures and field orientations, revealing Shubnikov de-Haas (SdH) oscillations (Figs. 4(d-i)). Strong SdH oscillations can be observed up to 20 K and 14 T after a smooth background subtraction, as can be seen in Fig. 4(d). The subsequent fast Fourier transform (FFT) on the SdH oscillations reveals that they consist of two frequencies Fα = 104 T and Fβ = 212 T (Fig. 4(e)). Despite the fact that Fβ is close to 2 $\times$ Fα, the effective mass fitting (inset of Fig. 4(e)) shows that the two frequencies originate from distinct Fermi surface cross-sectional areas, since the effective mass mβ is not twice that of mα. The small effective masses also imply that the underlying band dispersion is linear or close to linear.
[…]
The complex Fermi surface of InTaS2, as illustrated in Fig. 1(e) and Figs. 4(a-c), should lead to QOs of various frequencies, while the SdH oscillations up to 14 T in Figs. 4(d-g) only consists of two frequencies. To further incorporate larger Fermi pockets, we measured the SdH oscillations up to 44.8 T at different magnetic field orientations (Figs. 4(h-i), supplemental information Figs. S5, 6). The SdH oscillations (supplemental information Fig. S5(a,b)) show a much more complex spectrum with more frequencies, consistent with the complex Fermi surface of InTaS2. Most interestingly, we discovered QOs of frequencies 6 kT and 12 kT (Fig. 4(h)). By comparing the experimental results (symbols) to DFT calculations (dashed lines) in Fig. 4(i), we found that these frequencies are related to the oscillations from the Fermi pocket $\epsilon$ (see the right panel of Fig. 1(e)) and its second harmonic. Such $\epsilon$ pockets encode pinched points enforced by both the $\Gamma-M$ and $\Gamma-A$ KNLs. Moreover, the frequency of the Fermi pocket $\epsilon$ (symbols) is significantly lower than the expected value for a cylindrical Fermi pocket (solid curve in Fig. 4(i)), which indicates that it must be a closed pocket as in the DFT calculation. We notice that at $\theta$ $\sim$ 9-10, the fundamental oscillations from the Fermi pocket $\epsilon$ vanishes while the second harmonic oscillations persist. This implies that at this field orientation, the quantum oscillations from the Fermi pocket $\epsilon$ are close to the spin-zero effect [56]. Overall, through ARPES and quantum oscillation measurements, we directly observe the Fermi pockets related to the KNLs in InTaS2. The underlying quasiparticles could contribute to the properties of InTaS2, for instance, the superconductivity.
Discussions and Conclusions
The KNLs in noncentrosymmetric achiral little group have been clearly demonstrated in the exemplary case of the $P\bar{6}m2$ intercalated TMD superconductors consisting of InxTaS2 ($x$ = 1/2, 2/3, 1) and PbTaSe2, where In1/2TaS2 showcases the cleanest FS with a single KNL crossing the Fermi level, termed as the “ideal Kramers nodal line metal”. The spin-orbital texture in In1/2TaS2 directly observed via spin-resolved ARPES offers a natural explanation for the spin-valley polarization in the language of KNL, with a large SOC spin splitting up to around 250 meV. Furthermore in InTaS2, combined ARPES and angle-dependent QO studies point to the existence of the pinch points enforced by the KNLs, especially for the experimentally-probed closed $\epsilon$ pockets intersecting both the $\Gamma-M$ and $\Gamma-A$ KNLs. These pinch points are reminiscent of the 2D massless Dirac fermions on the surface of a 3D topological insulator, but reside on the intersection of the 3D FSs of an ideal KNL metal. Therefore, when the pinch points that harbor $n\pi$ ($n\in\mathbb{Z}$ and is an odd number) Berry phase are gapped by superconductivity, they produce nontrivial vortex spectra hosting chiral Majorana zero modes, which we elaborate in Supplementary Note 5 based on the model analysis of an ideal KNL. And in this work, both the KNLs with their associated pinch points and superconductivity are demonstrated in the InxTaS2 family. More importantly, our work establishes an entire large family of quantum materials as a platform for realizing and tuning KNLs – intercalated TMD compounds. Specifically, the In site can be populated with different concentrations of Tl, Pb, Bi, and Sn, the Ta site with Nb, and the S site with Se [32]. The intercalation not only produces the broken inversion symmetry – a central requirement for realizing KNLs, but can also be utilized to introduce correlated physics and electronic orders in the presence of KNLs, such as magnetic orders, charge-density waves, or superconductivity, a regime that has not been previously explored. Therefore, this material family offers promising ingredients for spin and valley transport, axionic quasiparticles, and topological superconductivity.
Going beyond material selection, two other directions would deserve further exploration. The first one is the interplay between dimensionality and unconventional superconductivity. Recently, unconventional nodal superconductivity has been reported in monolayer 1H-TaS2 [52]. In the exfoliable intercalated $P\bar{6}m2$ TMD family represented by indium intercalated 1H-TaS2 single crystals, the nature of the superconductivity deserves further investigation, especially upon approaching the few-layer limit. The second direction is strong electronic correlations. As analyzed in Supplementary Note 5, due to the Dirac physics of the KNL model, the AHE can be driven by Zeeman or exchange fields along the KNLs. However, the strength of the anomalous Hall conductivity is inversely proportional to the velocity of the KNL band. In the case of InxTaS2, the dispersive KNLs would give rise to an AHE that is overwhelmed by the ordinary Hall contributions. One way to increase the anomalous Hall signal is to induce flat KNLs at the Fermi level, which can be achieved in localized $f$-orbital electronic systems with one example in $P\bar{6}m2$ being UAuSi [57]. Such scenario may open an alternative avenue for topological phases and observables driven by strong correlations.
Acknowledgments
We thank Shiming Lei for discussion at the initial stage of the project. This work was mainly supported by the Department of Defense, Air Force Office of Scientific Research under Grant No. FA9550-21-1-0343. This research used resources of the Advanced Light Source and the Stanford Synchrotron Radiation Lightsource, both U.S. DOE Office of Science User Facilities under Contracts No. DE-AC02-05CH11231 and No. AC02-76SF00515, respectively. Part of the research also used resources of the National Synchrotron Light Source II, a U.S. Department of Energy (DOE) Office of Science User Facility operated for the DOE Office of Science by Brookhaven National Laboratory under Contract No. DE-SC0012704. The ARPES work at Rice University was also supported by the Gordon and Betty Moore Foundation’s EPiQS Initiative through grant No. GBMF9470. Y. Gao’s work at the National High Magnetic Field Laboratory was funded by the Gordon and Betty Moore Foundation’s EPIQS Initiative through ICAM-I2CAM, Grant GBMF5305. L. B. is supported by the US-DoE, BES program, through award DE-SC0002613. A portion of this work was performed at the National High Magnetic Field Laboratory, which is supported by National Science Foundation Cooperative Agreement No. DMR-2128556 and the State of Florida. Work at the University of California, Berkeley and Lawrence Berkeley National Laboratory was funded by the U.S. DOE, Office of Science, Office of Basic Science, Materials Sciences and Engineering Division under Contract No. DE-AC02-05CH11231 (Quantum Materials Program KC2202). We acknowledge support from the US National Science Foundation (NSF) Grant Number 2201516 under the Accelnet program of Office of International Science and Engineering (OISE). The KKR calculation work was supported by the project Quantum materials for applications in sustainable technologies (QM4ST), funded as project No. CZ.02.01.01/00/22_008/0004572 by P JAK, call Excellent Research. Research at UBC was undertaken thanks in part to funding from the Canada First Research Excellence Fund, Quantum Materials and Future Technologies Program. P.R.H. acknowledges funding from NSF grant no. DMR 2047193.
Methods
0.4 Crystal growth.
Single crystals of InxTaS2 ($x$ = 1/2, 2/3, 1) were grown by the chemical vapor transport (CVT) method. The polycrystalline precursor of InTaS2 was first prepared by a solid-state reaction. Indium powder, tantalum powder, and sulfur powder were mixed homogeneously with a mortar and pestle and sealed in quartz tubes under vacuum. The powder was heated up to 850C in 17 hours and kept at this temperature for 24 h before cooling down to room temperature. Approximately 3g of InTaS2 and the transport agent InCl3 (1mg/cm3 for InTaS2, 4mg/cm3 for In1/2TaS2 and In2/3TaS2) were put together in a quartz tube (200 mm length, inner diameter 16 mm). All treatments were carried out in an argon box, with an oxygen and water content of less than 0.5 p.p.m. The quartz tubes were sealed and put into a two-zone furnace. The hot end with the starting materials was kept at 1050C, and the cold end was kept at 1000C. Single crystals of InxTaS2 ($x$ = 1/2, 2/3, 1) can be found in the middle of the quartz tube after a growth of 7 days. The single crystals of different compositions can be further distinguished by XRDs due to the differences in the lattice constant $c$.
0.5 Transport and thermodynamic measurements.
The magnetization measurements were performed in a Quantum Design Magnetic Property Measurement System-3 (MPMS-3) magnetometer with a He-3 option. The magnetotransport measurements up to 14 T were measured in a Quantum Design Physical Property Measurement System (PPMS) dynacool system equipped with a dilution refrigerator option. The resistance was measured through the ETO option. The magnetotransport measurements on the same sample were measured at the National High Magnetic Field Laboratory (NHMFL) in Tallahassee in Cell-15, i.e, hybrid magnet, under magnetic field up to 44.8 T and temperature down to 0.35 K. The resistance was measured through a Lakeshore 370 AC resistance bridge.
0.6 ARPES measurements.
Angle-resolved photoemission spectroscopy (ARPES) and spin-ARPES measurements on In1/2TaS2 were collected at the Advanced Light Source, beamline 10.0.1.2 under $p$ polarized photons, equipped with a Scienta Omicron DA30L spectrometer. The In1/2TaS2 samples were cleaved in-situ with a base pressure better than $4\times 10^{-11}$ Torr and at a maintained temperature of 15 K. During spin-ARPES measurements, VLEED (very low-energy electron diffraction) detectors were used with the spin quantization axis fixed along the out-of-plane ($s_{z}$) direction in Fig. 3 and the spin polarization is calculated from
$P=\frac{1}{S}\frac{I_{\uparrow}-I_{\downarrow}}{I_{\uparrow}+I_{\downarrow}},$
where $S$ is the Sherman function. During the time of the experiment, the Sherman function took the value of 0.2. The corresponding spin-up and spin-down EDCs were measured up to the same acquisition time and normalized by the area using the background counts between 1.2 and 1.7 eV binding energies. The error bars of the spin polarization are calculated using the error propagation formula:
$\delta P=P\cdot\sqrt{\frac{(\sqrt{I_{\uparrow}})^{2}+(\sqrt{I_{\downarrow}})^{2}}{(I_{\uparrow}+I_{\downarrow})^{2}}+\frac{(\sqrt{I_{\uparrow}})^{2}+(\sqrt{I_{\downarrow}})^{2}}{(I_{\uparrow}-I_{\downarrow})^{2}}},$
where the uncertainty of the spin-resolved photoelectron counts takes the form of $\sqrt{I_{\uparrow}}$ and $\sqrt{I_{\downarrow}}$ assuming the Gaussian statistics of $I_{\uparrow}$ and $I_{\downarrow}$. In this calculation, the uncertainty from the Sherman function is not taken into consideration. The large error bar on one data point in Fig. 3(e) comes from the close values of spin up and spin down counts in the raw data. By definition, identical spin up and down counts give rise to a divergent uncertainty of the spin polarization. Additionally, the positions of $k_{3}$ and $k_{4}$ in Fig. 3(f) were actually taken on the opposite sides of the $\bar{K}-\bar{M}-\bar{K^{\prime}}$, closer to the $\bar{\Gamma}$ hexagonal pocket than the cut in Fig. 3(a). We instead showed the cut to be outside of the BZ just for illustration purpose of a complete FS image. The curvature of the momentum cut on the FS in Fig. 3(a) and (f) is omitted to avoid curved and distorted band dispersions in the 3D plot.
ARPES data on InTaS2 and PbTaSe2 were taken at the Stanford Synchrotron Radiation Lightsource, Beamline 5-2 and the Brookhaven National Lab, National Synchrotron Light Source II, Beamline 21-ID, respectively. Both are equipped with a DA30 electron analyzer with vertical slit and have linear horizontal, linear vertical, and circularly polarized photons available. Only the data measured by linear horizontal light were included in the main text and the supplemental information. Samples of both kinds were cleaved in-situ under a based pressure better than $3\times 10^{-11}$ Torr and temperatures below 30K. All ARPES measurements in the standard mode maintain an energy and angular resolution superior to 20 meV and 0.1, while spin-ARPES on In1/2TaS2 has the energy and angular resolution better than 50 meV and 1.
0.7 Ab-initio calculations
The Bloch spectral function and the one-step model ARPES calculations were carried out using the spin-polarized relativisitic Korringa-Kohn-Rostoker (SPR-KKR) package [58], under the full-potential fully-relativistic four component Dirac formalism. Exchange-correlation potential within the local spin density approximation by Vosko, Wilk, and Nusair [59] was used. To model the random vacancy on the indium site in InxTaS2 ($x=$ 1/2, 2/3), single-site coherent potential approximation was employed to obtain an auxiliary effective medium that reproduces the concentration averaged scattering properties [60, 58]. The KKR equations were solved with an angular momentum cutoff of $l_{max}$ = 4 to account for the occupied 4f orbitals of Ta and the needs for spectroscopy calculations. The ARPES calculations considered a semi-infinite surface model terminated by the In atoms under the experimental geometry and used the layer-KKR multiple scattering theory, together with coherent potential approximation [61, 62]. Therefore, such theory takes into account all factors such as light polarization, matrix element, final-state, surface, disorder, relativistic, and multiple scattering effects. Experimental crystal structures of In1/2TaS2 and In2/3TaS2 were utilized in the calculations [32, 33].
For the calculations on InTaS2, we used the Vienna Ab initio Simulation Package (VASP) [63] with the Perdew-Berke-Ernzerhof (PBE) exchange-correlation functional in the generalized-gradient approximation [64, 65] to perform the density functional theory (DFT) calculations [66]. The crystallographic data were obtained from the topological material database [67, 68, 69]. SOC is included in the first-principles calculations. The Wannier tight binding model was further obtained through the Wannier90 package [70], which accurately fits the DFT bands with the inclusion of SOC. From this Wannier model, we studied in detail the topological crossings in InTaS2. To better match with the ARPES data, we used a surface potential of -0.7 eV on the In $s$ orbitals for the In termination and +0.2 eV on the Ta $d$ orbitals for the S termination.
Data availability
The data that support the findings of this study are available from the corresponding authors upon reasonable request.
Author contributions
M.Y. and E.M. conceived the project. J.H. suggested the material family. Y.Gao synthesized the crystals and performed magnetotransport and magnetization measurements with the help from M.O. ,A.H., A.M., and L.B. under the supervision of E.M.. Y.Z., Y.Guo, Z.Y., and J.S.O. performed the ARPES experiments under the supervision of M.Y. and R.J.B, with the help from A.F., S.K.M, M.H., D.L., A.R., and E.V.. A.P. and Y.Z. performed the KKR calculations under the supervision of J.M.. X.G. carried out the wave-function based first-principles calculations, with the help from X.J.G., under the supervision of K.T.L.. P.H. provided theory support. Y.Z., Y.Gao, and A.P. contributed equally to this work. All authors contributed to the manuscript preparation.
Competing Interests
The authors declare no competing interests.